OpenStudy (anonymous):

A closed box with a square base is to be made from 300 cm2 of a certain material. Find the dimensions of the box with maximum volume.

OpenStudy (anonymous):

Let side of square be a a^2 = 300 Find the max for x using 1st derivative test

OpenStudy (anonymous):

could you be a little more specific in the procedure, please? thank you!

OpenStudy (anonymous):

Let x = length of side of base and y = height of box. Surface area \[A=2x^2+4xy\\ 300=2x^2+4xy\] Volume \[V=x^2y\] Find an expression for y in terms of x using the surface area equation: \[300=2x^2+4xy\\ \Rightarrow y=\frac{300-2x^2}{4x}\] Substitute this into the volume equation, then apply the first derivative test.

OpenStudy (anonymous):

thank you so much!!